# Shadow of Quantum Improved Regular Kerr Black Hole and parameter constrains with EHT observations

### v1

@@ III Null geodesics and Photon sphere
In the rotating case, the photon sphere cannot be fully determined by Eqs. (27) and (28) alone. The angular equation of motion (20) must also be considered, specifically requiring $\Theta\geq 0$, where
$\frac{\Theta(\theta)}{E^{2}}=\eta+\left(a^{2}-\frac{\xi^{2}}{\sin^{2}\theta}\right)\cos^{2}\theta\geq 0.$ (33)
Figure 4: Plot of $\Theta(\theta)/E^{2}$ as a function of $\xi$ and $\eta$. There are two types of $\theta$-motion modes. The left panel corresponds to $\xi^{2}\geq a^{2}$, while the right panel corresponds to $\xi^{2}<a^{2}$.
As shown in Fig. 4, the condition $\Theta\geq 0$ leads to two cases:
  • Case 1: When $\xi^{2}\geq a^{2}$ and $\eta\geq 0$, $\Theta(\theta)$ has two zero points, denoted $\theta_{\text{min}}$ and $\theta_{\text{max}}$, such that $\theta_{\text{min}}+\theta_{\text{max}}=\pi$. In this case, the photon oscillates between $\theta_{\text{min}}$ and $\theta_{\text{max}}$, allowing it to cross the equatorial plane. If $\eta=0$, the photon remains confined to the equatorial plane at $\theta=\pi/2$.
  • Case 2: When $\xi^{2}<a^{2}$ and $(a-|\xi|)^{2}+\eta\geq 0$, the behavior differs. Here, $\eta$ may take negative values, altering the $\theta$-motion. For $\eta>0$, the photon’s $\theta$-motion resembles the first case. However, for $\eta<0$, $\Theta(\theta)$ has four distinct zero points: $\theta_{1}<\theta_{2}<\theta_{3}<\theta_{4}$. In this scenario, the photon oscillates within either $(\theta_{1},\theta_{2})$ or $(\theta_{3},\theta_{4})$, and it cannot cross the equatorial plane.
To fully solve for the photon sphere, it is necessary to simultaneously satisfy the three conditions in Eqs. (27), (28), and (33). For photon sphere associated with shadow, these conditions typically reduce to $\eta\geq 0$ in most rotating black holes. In contrast, orbits inside the event horizon may require consideration of cases where $\eta$ is negative.
Once the photon sphere range is determined, by substituting $r_{\text{sphere}}$ into Eqs. (27) and (28), we can compute the impact parameters $\xi$ and $\eta$. The stability of these orbits is then assessed by analyzing $R^{\prime\prime}(r_{\text{sphere}})$, while the permissible range for the photon’s $\theta$-motion is determined by the condition $\Theta\geq 0$, as specified in Eq. (33).
IV Analysis of shadow-related functions
In this section, we will use the analytical expressions for the photon sphere obtained in Sec. III to further investigate the structure of the photon sphere. We will study the asymptotic behavior of the shadow-related functions in a more general case to gain a more intuitive understanding of the photon sphere and its impact on the black hole shadows. Specifically, we will solve for the photon sphere of the QIRK black hole to illustrate these effects.
We consider four functions related to the black hole shadow: $\eta(r)$, $\xi^{2}-a^{2}$, $R^{\prime\prime}(r)$, and $\xi(r)$. The functions $\eta(r)$ and $\xi^{2}-a^{2}$ govern the photon’s mode of motion in the $\theta$-direction, defining the allowed region for photon motion, as discussed in Sec. III. The function $R^{\prime\prime}(r)$ characterizes the stability of the photon sphere, indicating whether a photon sphere at a given radius is stable and observable. Furthermore, $\xi(r)$ directly correlates with the celestial transverse coordinate of the shadow in (42). The monotonicity of $\xi(r)$ within the allowable range of photon sphere radii influences the shape of the shadow. Notably, if $\xi(r)$ exhibits non-monotonic behavior, this may give rise to a ”cuspy-like” shadow structure Cunha:2017eoe .
By examining the expressions for $\xi$ and $\eta$ in (27) and (28), along with the second derivative of $R$ in (30), we can analyze the asymptotic behavior of these functions as $r\rightarrow\infty$, $r\rightarrow r_{+}$, and $r\rightarrow r_{0}$, where $r_{0}$ is the zero point of $\Delta^{\prime}(r)$. Assuming $\Delta(r)\rightarrow r^{n}$ as $r\rightarrow+\infty$, we obtain the following results:
  1. 1.
    As $r\rightarrow\infty$:
    $\xi(r)$ $\rightarrow\frac{n-4}{an}r^{2}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}-\infty,\quad\eta(r)\rightarrow\frac{-n^{2}+8n-16}{a^{2}n^{2}}r^{4}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}-\infty,$
    $R^{\prime\prime}(r)$ $\rightarrow\frac{8(4-n)}{n}r^{2}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}+\infty,\quad\xi^{2}-a^{2}\rightarrow\Big{(}\frac{n-4}{an}\Big{)}^{2}r^{4}-a^{2}\stackrel{{\scriptstyle n=2}}{{\longrightarrow}}+\infty.$
    Here, the asymptotic behavior of the four functions depends on the asymptotic form of $\Delta(r)$ as $r\rightarrow\infty$, particularly for QIRK black holes where $n=2$.
  2. 2.
    In the non-extremal case, $r_{-}\neq r_{+}\neq r_{0}$:
    1. (a)
      As $r\rightarrow r_{+}^{+}$:
      $\xi(r_{+})$ $=\frac{a^{2}+r_{+}^{2}}{a}>0,\quad\quad\quad\eta(r_{+})=-\frac{r_{+}^{4}}{a^{2}}<0,$
      $R^{\prime\prime}(r)$ $=8r_{+}^{2}>0,\quad\quad\quad\xi^{2}-a^{2}=2r_{+}^{2}+\frac{r_{+}^{4}}{a^{2}}>0.$
    2. (b)
      As $r\rightarrow r_{0}^{+}$:
      $\xi(r)$ $\rightarrow\frac{-4r_{0}\Delta(r_{0})}{a\Delta^{\prime}(r_{0})}\rightarrow+\infty,\quad\eta(r)\rightarrow\frac{16r_{0}^{2}\Delta(r_{0})(a^{2}-\Delta(r_{0}))}{a^{2}\Delta^{\prime}(r_{0})^{2}}\rightarrow-\infty,$
      $R^{\prime\prime}(r)$ $\rightarrow-\frac{16r_{0}^{2}\Delta(r_{0})\Delta^{\prime\prime}(r_{0})}{\Delta^{\prime}(r_{0})^{2}}\rightarrow+\infty,\quad\xi^{2}-a^{2}\rightarrow\frac{16r_{0}^{2}}{a^{2}}\Big{(}\frac{\Delta(r_{0})}{\Delta^{\prime}(r_{0})}\Big{)}^{2}-a^{2}\rightarrow+\infty.$
    In this case, the divergence of the four functions at the boundaries $r\rightarrow\infty$ and $r\rightarrow r_{0}$ makes them behave similarly to quadratic functions within this interval. Assuming all four functions are quadratic-like, these asymptotic behaviors allow us to deduce the opening direction of these curves, which provides an intuitive view of the photon sphere structure.
  3. 3.
    In the extremal case, $r_{-}=r_{+}=r_{0}$, when $r\rightarrow r_{0}$ and $\Delta^{\prime\prime}(r_{0})\neq 0$:
    $\xi(r_{0})$ $=\frac{a^{2}+r_{0}^{2}}{a}>0,\quad\eta(r_{0})=r_{0}^{2}\frac{8a^{2}-r_{0}^{2}\Delta^{\prime\prime}(r_{0})}{a^{2}\Delta^{\prime\prime}(r_{0})},$
    $R^{\prime\prime}(r_{0})$ $=0,\quad\quad\quad\xi^{2}-a^{2}=2r_{0}^{2}+\frac{r_{0}^{4}}{a^{2}}>0.$
    Here, the value of $\eta(r_{0})$ depends on the specific black hole model, and its sign determines whether the photon sphere can extend continuously inside the event horizon. This may subsequently affect the analysis of the black hole shadow. Further discussion on this will be provided in Sec. V.
In each case above, the first arrow or equality represents the general scenario, while the following expression is specific to the QIRK black hole. Moreover, by analyzing their monotonicity, we find that:
$R^{\prime\prime}(r)$ $=8r\left(\frac{2\Delta(r)\left(\Delta^{\prime}(r)-r\Delta^{\prime\prime}(r)\right)}{\Delta^{\prime}(r)^{2}}+r\right),$ (34)
$\xi^{\prime}(r)$ $=-\frac{2}{a}\left(\frac{2\Delta(r)\left(\Delta^{\prime}(r)-r\Delta^{\prime\prime}(r)\right)}{\Delta^{\prime}(r)^{2}}+r\right),$ (35)
$\eta^{\prime}(r)$ $=\frac{4r(4a^{2}-4\Delta(r)+r\Delta^{\prime}(r))}{a^{2}\Delta^{\prime}(r)}\left(\frac{2\Delta(r)\left(\Delta^{\prime}(r)-r\Delta^{\prime\prime}(r)\right)}{\Delta^{\prime}(r)^{2}}+r\right).$ (36)
These functions all share the same factor. When this factor equals zero at some $r\neq 0$, both $R^{\prime\prime}(r)$ and $\xi^{\prime}(r)$ will simultaneously vanish. However, $\eta^{\prime}(r)$ cannot be determined due to an additional term. This suggests that the stability of the photon sphere may change along with the monotonicity of $\xi(r)$, as can be observed in the extremal case.
Figure 5: $R^{\prime\prime}(r)$, $\eta(r)$, $\xi(r)$, and $\xi^{2}-a^{2}$ as functions of the photon sphere radius $r_{\text{sphere}}$. The vertical black dashed line represents the horizon. The top two panels show the QIRK BH in the non-extremal case and the Kerr BH with the corresponding rotation parameter $a$. The behavior of the four functions in both panels is highly similar, suggesting that the shadows of these black holes will exhibit comparable characteristics. The bottom two panels depict the extremal QIRK BH and the extremal Kerr BH, both of which display analogous behavior. The behavior of extremal QIRK black holes is discussed in further detail in Sec. V.
In Fig. 5, we have plotted the functions $R^{\prime\prime}(r)$, $\eta(r)$, $\xi(r)$, and $\xi(r)^{2}-a^{2}$ for both the QIRK black hole and the Kerr black hole, which align with the expected asymptotic behavior. Fig. 5 also provides a clear visualization of the photon sphere range and their properties.
In the non-extremal case, the behavior of these four shadow-related functions for the QIRK black hole closely resembles that of the Kerr black hole. Specifically, the range of $r_{\text{sphere}}$ lies outside the horizon, where $R^{\prime\prime}(r)>0$, indicating that they are unstable and directly contribute to the formation of the black hole shadow. Given the similar behavior of these functions in both the QIRK and Kerr black holes, we can infer that the shadow of the QIRK black hole will closely resemble that of the Kerr black hole. This similarity will be further explored in Sec. V.
In the extremal case, however, the functions $\eta(r)$, $\xi^{2}-a^{2}$, $R^{\prime\prime}(r)$, and $\xi(r)$ do not diverge at $r_{0}$ any more. Analyzing the photon region, we observe that $\eta(r_{0})>0$, implying that the region where $\eta(r)\geq 0$, corresponding to the photon region, extends continuously inside the horizon. Moreover, we find $R^{\prime\prime}(r_{0})=0$, indicating a transition in the stability of the photon sphere: while the photon sphere outside the horizon remains unstable, the one inside the horizon is stable. It is important to note that this stable photon sphere does not contribute to the shadow. As illustrated in Fig. 5, this structure closely resembles that of the extremal Kerr black hole.
If we consider the photon sphere located at the horizon, the equations (25) and (26) simplify to:
$\frac{R(r_{0})}{E^{2}}$ $=\left(a^{2}-a\xi+r_{0}^{2}\right)^{2}=0,$ (37)
$\frac{R^{\prime}(r_{0})}{E^{2}}$ $=4r_{0}\left(a^{2}-a\xi+r_{0}^{2}\right)=0.$ (38)
There is no restriction on $\eta$. This means that as long as the photon’s impact parameter $\xi$ satisfies Eq. (37) and Eq. (38), the photon can move in spherical orbits at the horizon.
V Shadow
In general, photons emitted from a light source are deflected by a black hole due to gravitational lensing. These photons can be categorized into three types: scattered, bounded, and absorbed. Scattered photons reach distant observers, while bounded and absorbed photons orbit near the black hole and fall into it, respectively. This creates a dark region on the observer’s plane, known as the black hole shadow. The boundary between the light and dark regions is defined by the orbits of these bounded photons.
For an observer at the position $(r_{obs},\theta_{0})$, in the far exterior region of the black hole, The celestial coordinates $(\alpha,\beta)$ are obtained by projecting the velocity of the photons observed by the observer onto the horizontal and vertical coordinates of the observation plane 1973blho.conf..215B ; Chandrasekhar_1985 , i.e.,
$\alpha=-r_{obs}\frac{p^{\varphi}}{p^{t}},\quad\beta=r_{obs}\frac{p^{\theta}}{p^{t}},$ (39)
where $p^{\mu}$ is the photon four-momentum measured. For an observer sitting in the asymptotically flat region $r_{obs}\rightarrow\infty$, substitute into the geodesic equation, celestial coordinates can be simplified as
$\alpha$ $=$ $\lim_{r_{obs}\to\infty}\Big{(}-r^{2}_{obs}\sin\theta_{0}\frac{\mathrm{d}\varphi}{\mathrm{d}r}\bigg{|}_{\theta=\theta_{0}}\Big{)}=-\frac{\xi}{\sin\theta_{0}}\,,$ (40)
$\beta$ $=$ $\lim_{r_{obs}\to\infty}\Big{(}r^{2}_{obs}\frac{\mathrm{d}\theta}{\mathrm{d}r}\bigg{|}_{\theta=\theta_{0}}\Big{)}=\pm\sqrt{\eta+a^{2}\cos^{2}\theta_{0}-\xi^{2}\cot^{2}\theta_{0}}\,,$ (41)
where $r_{obs}$ denotes the distance between the observer and the black hole, $\theta_{0}$ represents the inclination angle between the observer’s line of sight and the rotational axis of the rotating black hole. When the observer is on the equatorial plane of the black hole with the inclination angle $\theta_{0}=\pi/2$, we obtain
$\alpha$ $=-\xi\,,$ (42)
$\beta$ =±$=\pm\sqrt{\eta}\,.$ (43)
In this situation, the celestial coordinates reduce to the impact parameters of the photons, which are shadow-related functions discussed in Sec. IV.
V.1 Non-Extremal Case
Figure 6: The Quantum Improved Regular Kerr black hole shadows with varying parameters $a$, $p$ and $\widetilde{\omega}$.
In the non-extremal case $(\widetilde{\omega}_{m}<\widetilde{\omega}<\widetilde{\omega}_{c})$, Fig. 6 shows the shadows of QIRK black hole that for fixed rotation $a$ and $p$. As $\widetilde{\omega}$ increases from $\widetilde{\omega}_{m}$ to near the critical value $\widetilde{\omega}_{c}$, we find that the leftmost point of intersection of the shadow with the $\alpha$-axis gradually moves inward, deviating from both the Kerr case and the standard circular shadow. However, the rest of the shadow remains mostly unchanged. Additionally, for fixed $a$ and $\widetilde{\omega}$, increasing $p$ causes the shadows to converge toward those of the Kerr black hole with the same rotation $a$. This suggests that $p=4$ represents the maximum deviation from the Kerr shadow. When $p$ and $\widetilde{\omega}$ are fixed, varying $a$ produces effects similar to those seen in general rotating black holes – larger values of $a$ cause the shadow to deviate more from the standard circular shape.
Overall, in the non-extremal case, smaller values of $p$ and larger values of $\widetilde{\omega}$ lead to greater deviations from the Kerr shadow, particularly where the leftmost point of the shadow intersects the $\alpha$-axis, while the rest of the shadow remains relatively unchanged regardless of $p$ and $\widetilde{\omega}$.
V.2 Extremal Case
In the extremal case $(\widetilde{\omega}=\widetilde{\omega}_{c})$, the inner and outer horizons coincide. For Kerr black holes, achieving extremality requires $a=1$. However, the QIRK black hole can reach the extremal case for any allowable $a$ by setting $\widetilde{\omega}=\widetilde{\omega}_{c}$, as indicated by the black curve in Fig. 3. This extremal case presents notable differences from the non-extremal scenario. Fig. 5 illustrates the shadow-related functions for an extremal QIRK black hole with $p=4$ and $a=0.9$, contrasted with the extremal Kerr black hole. In this extremal configuration, none of the four functions diverge near the horizon.
When considering photon sphere that contribute to the shadow, only cases with $\eta(r)>0$ outside the horizon are relevant, as the photon sphere within the horizon is stable and thus not observable. Consequently, the range of $r_{\text{sphere}}$ should not extend inside the horizon. Notably, however, we observe that $\eta(r_{+})>0$, meaning that when plotting the shadow’s curve with $r_{\text{sphere}}$ as a parameter, the resulting curve will not be closed, since its vertical coordinate $\beta(r_{+})=\sqrt{\eta(r_{+})}>0$. This behavior stems from the unique properties of the photon sphere at the horizon. As discussed in Sec. IV, the photon sphere at the horizon imposes a constraint on $\xi$, while $\eta$ can vary freely, forming a straight line parallel to the $\beta$-axis in the observation plane. This line intersects the open curve, ultimately completing it as a closed loop.
This behavior, however, does not hold across all extremal cases. Fig. 7 shows that for fixed $p$, there exists a critical value $a_{\text{cri}}$, above which the region with $\eta(r)>0$ extends into the horizon. For $a<a_{\text{cri}}$, the scenario resembles that of the non-extremal case, as suggested by our discussion of the extremal case for $\eta(r_{0})$ in Sec. IV. The specific values of $a_{\text{cri}}$ for each $p$ are listed in Table 1.
In Fig. 8, the first panel provides a clear visualization of the discussion in Sec. IV. The dashed lines in all four panels mark the formation of photon sphere at the horizon. Shadows only appear on these dashed lines when $a>a_{\text{cri}}$, not when $a<a_{\text{cri}}$, which aligns with the behavior of the extremal Kerr case. Additionally, the trends in shadow shape and size are consistent with those observed in the non-extremal case: smaller $p$ and larger $a$ result in shadows that deviate further from the standard circular form.
Figure 7: In the extremal case, the relationship between $\eta(r)$ and the horizon with fixed $p$. Different curves correspond to $\eta(r)$ for various rotation parameters $a$, where the vertical lines indicate the positions of the horizon for each corresponding $a$. There exists a critical rotation parameter $a_{\text{cri}}$. For $a>a_{\text{cri}}$, the region where $\eta(r)>0$ extends into the horizon, while for $a<a_{\text{cri}}$, this region remains outside the horizon.
Figure 8: Shadows of the extremal QIRK black hole with varying parameters $a$ and $p$. The solid curves represent the shadow formed by the photon sphere outside the horizon, while the dashed lines correspond to the photon sphere located at the horizon. When $a>a_{\text{cri}}$, both dashed lines and solid curves appear. Conversely, for $a<a_{\text{cri}}$, only solid curves are present.
Overall, the effect of the parameter $p$ and the quantum correction $\widetilde{\omega}$, introduced by the QIRK metric, on black hole shadows follows the same trend for both non-extremal and extremal cases. Specifically, smaller $p$ and larger $\widetilde{\omega}$ within the allowed range $(\widetilde{\omega}_{m}<\widetilde{\omega}\leq\widetilde{\omega}_{c})$ lead to greater deviations from Kerr black holes. However, in terms of the magnitude of this effect on the shadows, when the rotation parameter $a$ is fixed and $p$ and $\widetilde{\omega}$ are varied across their full ranges, the change in the shadows is relatively small compared to many other rotating black holes.
Moreover, the change in the shadow of the QIRK black hole is primarily concentrated at the left intersection of the shadow with the $\alpha$-axis, similar to other rotating black holes based on asymptotic safe gravity Held:2019xde . In other words, even when the quantum effect is significant, its impact on the black hole shadow appears minimized, making it less prominent.
Another noteworthy point is the indistinguishability of extremal QIRK black holes. In the case of extremal Kerr black holes, the strictly straight line on the left side of the shadow makes them distinguishable from their non-extremal counterparts. However, for QIRK black holes, when $a<a_{\text{cri}}$, the shadow of the extremal QIRK black hole does not exhibit a strictly straight line at the leftmost edge, which is similar to the non-extremal case. As a result, it becomes indistinguishable from the non-extremal case. Only when $a>a_{\text{cri}}$ does the extremal QIRK black hole displays a shadow resembling that of the extremal Kerr black hole. Therefore, it is impossible to determine whether a QIRK black hole is in the extremal state based solely on the shape of its shadow.
VI Constraints from EHT Observations
In this section, we utilize observational results from the Event Horizon Telescope (EHT) for the supermassive black holes M87* and Sgr A* to constrain the parameters of the QIRK black hole model. By doing so, we aim to refine the QIRK black hole model, aligning it more closely with astronomical observations. To achieve this, we analyze three shadow observables: the circularity deviation $\Delta C$, the shadow angular diameter $\theta_{d}$, and the Schwarzschild deviation $\delta$, each constrained by EHT data.
The black hole shadow boundary is characterized as a one-dimensional closed curve described by radial and angular coordinates $(R(\varphi),\varphi)$ in a polar coordinate system centered at $(\alpha_{C},\beta_{C})$. The shadow’s average radius $\bar{R}$ is defined as follows johannsen_testing_2010 :

### v2

@@ III Null geodesics
The $\theta$-equation is identical to that in Kerr. It has been already discussed in great detail in Ref. Gralla:2019ceu , and we directly adopt the results here.
Table 2: Classification of photon trajectories based on the value of $r_{\max}$. Here, $r_{+}$ denotes the outer event horizon of the black hole.
Region Black Hole Case Compact Object Case
I $r_{\max}>r_{+}$ $r_{\max}>0$
II $0<r_{\max}<r_{+}$ $r_{\max}>0$
III $0<r_{\max}<r_{+}$ $r_{\max}>0$
IV $r_{\max}<0$ $r_{\max}<0$
V $r_{\max}$ is complex, i.e., light travels from $\infty$- ∞ to +$\infty$+ ∞ Same as black hole case
Fig. 4 shows the $(\xi,\eta)$ parameter plot for photon trajectories in both the black hole and compact object cases. Red curves denote unstable photon spheres, while blue curves denote stable photon spheres. These curves divide the parameter space into five regions, classified by the largest real root $r_{\max}$ of the radial potential $R(r)$. Since we focus on light received or emitted by an external observer at large $r$, only these regions are relevant, and the detailed classification is presented in Table 2. The red-shaded region, as in the Kerr case, is excluded by the $\theta$-motion. As stated earlier, QIRK introduces an additional family of photon spheres compared to the Kerr case Gralla:2019ceu , which appears as the straight line $\eta=0$ in the figure, and is unstable when $\lvert\xi\rvert<a$, and stable when $\lvert\xi\rvert>a$.
Figure 4: The $\xi,\eta$ parameter plot for a black hole ($a=0.8,p=4,\widetilde{\omega}=0.7$, left) and a compact object ($a=1.05,p=4,\widetilde{\omega}=0.1$, right). Here $\xi$ and $\eta$ are the usual impact parameters (see Eq. 23). The red curves represent unstable photon spheres. The blue curves represent stable photon spheres. Compared with Kerr, QIRK admits an extra photon sphere at $\eta=0,r=0$. In this case, $\lvert\xi\rvert<a$ is unstable, $\lvert\xi\rvert>a$ is stable
IV QIRK Black hole image with a thin accretion disk
Observations of black hole images offer a direct way to probe extreme gravitational fields, test gravitational theories, and reveal the fundamental properties of black holes. However, interpreting these images requires separating the effects of spacetime geometry from the complex radiation of the surrounding accreting matter. In this context, modeling black hole images with a thin accretion disk is useful because it simplifies the distribution and emission properties of the disk, the thin-disk model highlights fine features caused by strong gravitational lensing, such as the inner shadow and photon ring. The QIRK black hole, constructed within the asymptotic safety framework, removes the ring singularity of the classical Kerr solution and introduces quantum corrections that alter the spacetime geometry. Consequently, the detailed shape, size, and location of the photon ring in a QIRK image illuminated by a thin disk directly encode information about these quantum corrections. This provides a basis for testing quantum gravity models, distinguishing QIRK black holes from Kerr black holes, and constraining quantum gravity parameters through observation.
Since the setup and methods have been detailed in previous works Hou:2022eev ; Meng:2025ivb , we provide only a brief summary here. We place a distant zero-angular-momentum-observer (ZAMO) at ($t_{o}=0$, $r_{o}$, $\theta_{o}$, $\phi_{o}=0$) and define the local orthonormal tetrad $e_{(\mu)}$ as
[…]
The photon’s four-momentum $k_{\mu}$ can be projected onto this tetrad as
$p_{(\mu)}=k_{\nu}\,e^{\nu}{}_{(\mu)},\quad p^{(\mu)}=\eta^{(\mu)(\sigma)}\,e^{\nu}{}_{(\sigma)}\,k_{\nu}\,,$ (35)
then we can define the so-called celestial coordinates $\Theta$ and $\Psi$ in the ZAMO frame to label each light ray, i.e.,
$\cos\Theta$ $=\frac{p^{(1)}}{p^{(0)}},\quad\tan\Psi=\frac{p^{(3)}}{p^{(2)}}.$ (36)
Using a stereographic projection, one can map these angles onto a two-dimensional screen via
[…]
Given the screen coordinates $x$ and $y$, we first compute the photon’s four momentum using (35), (36) and (37). Then, by ray tracing, we obtain the photon’s trajectory, from which we extract the redshift and intensity information carried by the photon.
We model the illumination source as an optically and geometrically thin accretion disk in the equatorial plane ($\theta=\pi/2$). Tracing light rays backward from the observer, they can intersect the equatorial accretion disk multiple times. At each intersection, the ray collects additional emission and contributes to the detected intensity. Thus, it is essential to identify the radius $r_{n}(x,y)$ of the $n$-th intersection during ray tracing. In particular, $n=1$ produces the direct emission image, $n=2$ yields the lensed ring image, and $n\geq 3$ corresponds to the photon ring image.
The observed intensity $I_{\text{obs}}$ at a frequency $\nu_{0}$ is obtained by summing the contributions from each intersection of the light ray with the disk. Each contribution is the disk-emitted specific intensity $I_{\mathrm{em}}$ multiplied by the redshift factor $g_{n}=\nu_{\mathrm{em}}/\nu_{0}$ Hou:2022eev :
$I_{\text{obs}}(\nu_{0},\alpha,\beta)=\sum_{n=1}^{N_{\text{max}}}g_{n}^{3}I_{\text{em}}(\nu_{\text{em}},r_{n}),$ (38)
where $I_{\text{em}}(r_{n})$ represents the specific intensity emitted by the accretion disk material in its rest frame at radius $r_{n}$. In this work, we adopt a phenomenological emissivity model commonly used in black hole imaging simulations Chael:2021rjo . Specifically, the emissivity is taken to follow a simple radially decreasing profile, and is independent of the emitted frequency over the narrow band of observation for simplicity. A common functional form for such a profile, as used to model M87* and Sgr A*, involves an exponential decay with a logarithmic radial dependence Chael:2021rjo ; Hou:2022eev
$I_{em}(r)=\exp\left[-\frac{1}{2}\left(\log\frac{r}{r_{h}}\right)^{2}-2\left(\log\frac{r}{r_{h}}\right)\right].$ (39)
The redshift factor $g_{n}$ depends on the photon’s four-momentum $k_{\nu}$ and the four velocity of the emitting material in the disk $u^{\mu}_{\text{disk}}$:
$g_{n}=\frac{k_{\mu}u^{\mu}_{\text{disk}}}{k_{\nu}u^{\nu}_{\text{obs}}},$ (40)
where $u^{\mu}_{\text{obs}}$ is the observer’s four velocity. To determine the four-velocity of the emitting material $u^{\mu}_{\text{disk}}$, we adopt a standard model for the accretion flow model. Material outside the innermost stable circular orbit (ISCO) follows stable circular orbits in the equatorial plane. Upon reaching the ISCO, the material plunges inward toward the event horizon, following geodesics that preserve its energy and angular momentum at the ISCO. Fig. 5 shows the radius of the ISCO as a function of the quantum correction parameter $\widetilde{\omega}$. As $\widetilde{\omega}$ increases, the ISCO radius shrinks for both prograde and retrograde orbits. However, these changes are small, with only the prograde orbits at high spin values exhibiting relatively noticeable shifts.
Figure 5: The radius of the ISCO as a function of the quantum correction parameter $\widetilde{\omega}$. Dashed curves represent retrograde orbits, while solid curves represent prograde orbits. The magenta, blue, black, and red curves correspond to spin parameters $a=0.1$, $0.5$, $0.8$, and $0.9$, respectively. Additionally, the black dash-dotted curves mark the ISCO radius at $\widetilde{\omega}=\widetilde{\omega}_{c}(a)$ for each spin (the extremal black hole), and the orange dash-dotted curves mark the ISCO radius at $\widetilde{\omega}=\widetilde{\omega}_{m}(a)$ (the threshold for avoiding closed timelike curves).
The effects of observer inclination $\theta_{o}$ and black hole spin $a$ have been studied previously in Hou:2022eev , following the discussion there, we place the observer at a distance $r_{o}=100$, with inclination $\theta_{o}=80^{\circ}$, and investigate the impact of quantum corrections on the image at different spin values. Moreover, because the parameter $p$ has only a minor effect and larger $p$ complicates the analysis, we set $p=4$ and focus on varying the quantum correction parameter $\widetilde{\omega}$.
(a) Kerr $a=0.1$
(b) $a=0.1$ $p=4$ $\widetilde{\omega}=0.7$
(c) $a=0.1$ $p=4$ $\widetilde{\omega}=1.9$
(d) $a=0.1$ $p=4$ $\widetilde{\omega}=4$
(e) Kerr $a=0.8$
(f) $a=0.8$ $p=4$ $\widetilde{\omega}=0.7$
(g) Kerr $a=0.9$
(h) $a=0.9$ $p=4$ $\widetilde{\omega}=0.3$
Figure 6: Images of the QIRK black hole illuminated by prograde accretion flows for selected model parameters.
(a) $a=0.1$
(b) $a=0.8$
(c) $a=0.9$
(d) $a=0.1$
(e) $a=0.8$
(f) $a=0.9$
Figure 7: The observed intensity distribution along x-axis (upper panel) and y-axis (bottom panel) for selected parameters.
Figure 6 shows that the images of QIRK black holes illuminated by a thin accretion disk exhibit several notable features. The central dark region, known as the inner shadow, corresponds to the direct image of the event horizon. Outside this region, the brightest ring structure is the photon ring, formed by unstable spherical photon orbits. It can be observed that as the spin parameter $a$ increases, the photon ring becomes more distorted and deviates further from a circle. The up-down asymmetry of the image is largely determined by the observer’s inclination angle, while the left-right asymmetry arises mainly from the rotation of the black hole and the orbital motion of the accreting matter. Our results show that the quantum correction parameter $\widetilde{\omega}$ produces only modest changes in the size and shape of the inner shadow and photon ring. However, as $\widetilde{\omega}$ increases, the most pronounced effect is a systematic decrease in overall image brightness, as seen in Fig. 7. This brightness reduction is present for all spin values $a$. Taken together with the modifications to the ISCO shown in Fig. 5, these findings summarize the overall impact of quantum corrections on both the ISCO and the null geodesics outside the event horizon.
(a) Kerr $a=0.1$
(b) $a=0.1$ $p=4$ $\widetilde{\omega}=0.7$
(c) $a=0.1$ $p=4$ $\widetilde{\omega}=1.9$
(d) $a=0.1$ $p=4$ $\widetilde{\omega}=4$
(e) Kerr $a=0.8$
(f) $a=0.8$ $p=4$ $\widetilde{\omega}=0.7$
(g) Kerr $a=0.9$
(h) $a=0.9$ $p=4$ $\widetilde{\omega}=0.3$
Figure 8: The redshift factor distributions of the direct images of the accretion disk. The red and blue indicate redshift and blueshift, respectively. And the red dot in each plot indicates the maximal blueshift point.
(a) Kerr $a=0.1$
(b) $a=0.1$ $p=4$ $\widetilde{\omega}=0.7$
(c) $a=0.1$ $p=4$ $\widetilde{\omega}=1.9$
(d) $a=0.1$ $p=4$ $\widetilde{\omega}=4$
(e) Kerr $a=0.8$
(f) $a=0.8$ $p=4$ $\widetilde{\omega}=0.7$
(g) Kerr $a=0.9$
(h) $a=0.9$ $p=4$ $\widetilde{\omega}=0.3$
Figure 9: The redshift factors distribution of the lensed images of the accretion disk. The red and blue indicate redshift and blueshift, respectively. And the border of the black region in each plot gives the lensed images of event horizon.
Fig. 8 and 9 display the redshift distributions of the direct image ($n=1$) and the lensed image ($n=2$) of the accretion disk. The plots show several features. The redshift includes contributions from both gravitational redshift and Doppler shift. On the left side of each plot, a noticeable blueshift appears due to the prograde motion of the accretion flow. If the flow were retrograde, the pattern would change significantly. As a result, the overall redshift map is dominated by gravitational redshift as a background, with Doppler effects providing the main variation. Under the current observational setup and parameter choices, quantum corrections cause only minor changes to this macroscopic appearance, including the magnitude of the redshift factor $g_{n}$. This means that distinguishing QIRK black holes with different $\widetilde{\omega}$, or separating QIRK from Kerr black hole based solely on the large scale morphology of the redshift map, is likely to be difficult. A more reliable way to distinguish them will require a precise quantitative analysis of geometric observables, which we discuss in the next section.
V Constraints from EHT Observations
The previous analysis of QIRK black hole images suggests that quantum corrections induce only subtle modifications to the main features, manifesting primarily as minor shifts in intensity or geometry. Although these qualitative findings are very useful, a more rigorous quantitative analysis is required for precise characterization. Certain observables that describe the geometric features of the shadow provide this capability. In this section, we analyze three geometric observables of the black hole shadow: the circularity deviation $\Delta C$, the shadow angular diameter $\theta_{d}$, and the Schwarzschild deviation $\delta$. Assuming that the supermassive black holes M87* and Sgr A* are described by the QIRK model, we use Event Horizon Telescope (EHT) observations to constrain the QIRK parameters. Through this procedure, we aim to investigate the quantum corrections introduced by QIRK in greater detail, and align the QIRK model more closely with astronomical data.
The black hole shadow boundary is characterized as a one-dimensional closed curve described by radial and angular coordinates $(R(\varphi),\varphi)$ in a polar coordinate system centered at $(x_{C},y_{C})$. The shadow’s average radius $\bar{R}$ is defined as follows johannsen_testing_2010 :